I am running out of time to finish my summer work, and yet I still do not know how to solve these problems:


"What values are excluded from the domain of the function?"


f(x)= 1/x-1 (one over x minus one)


another is f(x)= 3x+2




A friend of mine who is doing the same work told me to plug 0 in for x, but I am still not sure.

Answer by jake_6233(74)

f(x)= 1/x-1 (one over x minus one)
for simplicity think of domain as x values and range as y values. The question asks what values of x are not allowed. Remembering that division by 0 is not allowed. You want all values of x such that x-1=0 (these are not permitted) so solving our equation we get x=1.
1 is excluded from the domain of f(x)=1/x-1
f(x)= 3x+2 Remember that any value of x you plug in will yield a valid value for y so there are no numbers excluded from the domain of this function